OpenStudy (anonymous):

Completely Lost Please Help!! If f(x) = sin x and g(x) = cos x determine the following. f(4n+1)(x) g(4n+1)(x) f(4n+2)(x) g(4n+2)(x) f(4n+3)(x) g(4n+3)(x) f(4n+4)(x) g(4n+4)(x) Now, use your answers to calculate F(157)(x) when F(x) = 4 sin x − 3 cos x.

OpenStudy (jamesj):

f' = cos f'' = -sin f''' = -cos f'''' = sin = f Likewise for g, there is a cycle four long. Hence \[ F^{(157)}(x) = F^{(4 \cdot 39 + 1)}(x) = F'(x) \]

OpenStudy (jamesj):

See what to do now?

OpenStudy (jamesj):

Talk to me

OpenStudy (anonymous):

so 4(cos)+1 4(-sin)+1) and so on

OpenStudy (jamesj):

What that means is the (4n+1)th derivative. E.g., \[ f'''' =f ^{(4)}, \ \ \ f''''' = f^{(5)} = f^{(4 \cdot 1 + 1)} \]

OpenStudy (jamesj):

\[ f' = f^{(1)}, \ \ f'' = f^{(2)}, \ \ f''' = f^{(3)}, \ ... \]

OpenStudy (anonymous):

oh so what is n?

OpenStudy (jamesj):

n is an integer. Hence if n = 0 4n + 1 if n = 1, 4n + 1 = 5 if n = 2, 4n + 1 = 9 etc. The idea is the 1st derivative is equal to the 5th derivative is equal to the ninth derivative etc. I.e., \[ f' = f^{(1)} = f^{(5)} = f^{(9)} = ... \] Likewise for 4n + 2: \[ f'' = f^{(2)} = f^{(6)} = f^{(10)} = ... \] and for 4n +3 and 4n + 4

OpenStudy (jamesj):

*correction: if n = 0, then 4n + 1 = 1

OpenStudy (anonymous):

so im going to have to find 9 derivatives? or more?

OpenStudy (jamesj):

No, the whole idea is you see the pattern so you only need to calculate 4 derivatives. For example, with sin f = sin f' = cos f'' = -sin f''' = -cos f'''' = sin = f I.e., the 4th derivative equals the original function. Hence the fifth derivative \[ f''''' = f^{(5)} \] is equal to the derivative of the 4th derivative. And the fourth derivative is the original function. Hence the 5th derivative is just the first derivative! \[ f^{(5)} = (f^{(4)})' = f' = f^{(1)} \]

OpenStudy (jamesj):

Likewise the 6th derivative is the second derivative, because \[ f^{(6)} = (f^{(5)})' = (f')' = f^{(2)} \]

OpenStudy (anonymous):

oh gotcha, so the answer will be one of the four sin cos, -sin or -cos since it just has x beside it and you plug in them for x?

OpenStudy (jamesj):

sort of. But can you see now that \[ f^{(1)} = f^{(5)} = f^{(9)} = f^{(13)} = ... \] i.e., \[ f^{(4n+1)} \] is the same for integers \( n \geq 0 \).

OpenStudy (anonymous):

yes i see, so sin(x) then would be for 4n+1

OpenStudy (jamesj):

This is what you have to show. And then show that \( f^{(2n+2)} \) is the same for all integers n. And show that \( f^{(2n+3)} \) is the same for all integers n And show that \( f^{(2n+4)} \) is the same for all integers n. === Then show the same relationships for g(x) = cos x

OpenStudy (jamesj):

*correction: 4n + 1, 4n + 2, 4n + 3, 4n + 4. Not 2n + ...

OpenStudy (anonymous):

0_o so since f(x) = sin x = = sin^2(x) for 4n+1)

OpenStudy (jamesj):

No. Let's go back to this: f = sin f' = cos f'' = -sin f''' = -cos f'''' = sin = f What is the 5th derivative?

OpenStudy (anonymous):

cos? or si

OpenStudy (jamesj):

remember that \( f^{(n)} \) means the nth derivative of f. It does NOT mean the nth power of f.

OpenStudy (jamesj):

Yes. The fifth derivative of f is written, in notation as \[ f^{(5)} \] That's just a definition of a piece of notation in mathematics. Now, by definition of the 5th derivative in itself, it is the derivative of the fourth derivative: i.e., \[ f^{(5)} = (f^{(4)})' \] Now, what is the fourth derivative of f equal to?

OpenStudy (anonymous):

sin

OpenStudy (anonymous):

sin

OpenStudy (jamesj):

right. Hence \[ f^{(5)} = (sin)' = cos \]

OpenStudy (jamesj):

that is, the 5th derivative is equal to the first derivative. In notation: \[ f''''' = f' \] or in the new compact notation of this question: \[ f^{(5)} = f^{(1)} \] Again, this does NOT mean the 5th power of f; it's the fifth derivative. This second formula is just a way of writing the first formula.

OpenStudy (jamesj):

So the question is asking you to prove several things. The first thing it asking you to prove is that \[ f^{(4n+1)} \] is the same for all integers \( n \geq 0 \). I.e., \[ f^{(1)} = f^{(5)} = f^{(9)} = f^{(13)} = ... \] Make sense?

OpenStudy (anonymous):

sobut isnt it asking for more than just f(1), f95) and f(9)?

OpenStudy (jamesj):

Yes. It's asking you to establish 8 formulae, at the top of your question: f(4n+1)(x) g(4n+1)(x) f(4n+2)(x) g(4n+2)(x) f(4n+3)(x) g(4n+3)(x) f(4n+4)(x) g(4n+4)(x) Where the first number here in brackets means the (4n+j)th derivative

OpenStudy (jamesj):

Then use that for the second part of the question as I indicated in my first answer. Anyway. Play around with this for a bit. I need to help someone else for a while.

OpenStudy (anonymous):

do you know what the anser to the very last one is i have all 8 but i need the last one